Find an equation of a line with slope -3 and through (2,4).
Answer by checkley71(8403)

Y=mX+b
4=-3*2+b
4=-6+b
4+6=b
10=b THUS THE EQUATION OF THIS LINE WITH A SLOPE OF -3 AND THROUGH (2,4) IS:
Y=-3X+10